# Hybrid Beamforming with Widely-spaced-array for Multi-user Cross-Near-and-Far-Field Communications

### v1

@@ II-B2 Sub-connected MU-WSA
In sub-connected WSA, the $L_{t}$ RF chains are divided uniformly into $K$ groups, each group connected to one subarray. Therefore, the number of phase shifters in the sub-connected system can be reduced by a factor of $K^{2}$ compared to the fully-connected system. As a result, the analog precoder has a block-diagonal format as
$\mathbf{F}_{\mathrm{RF}}=\begin{bmatrix}\mathbf{F}_{1}&0&\ldots&0\\ 0&\mathbf{F}_{2}&\ldots&0\\ \ldots&\ldots&\ldots&\ldots\\ 0&\ldots&0&\mathbf{F}_{K}\end{bmatrix},$ (6)
where $\mathbf{F}_{j}\in\mathbb{C}^{\frac{N_{t}}{k}\times\frac{L_{t}}{k}}$ represents the analog precoder for each subarray, and should satisfy $\left|\mathbf{F}_{j}(m,n)\right|=\frac{1}{\sqrt{N_{t}/K}},\forall j$.
@@ II-C Problem Formulation
Based on the system model of the MU-WSA architecture in both FC and SC systems, the beamformers can be designed to maximize the sum SE. Moreover, the subarray spacing $d_{s}$ and the number of subarrays $K$ need to be investigated since they have a significant impact on the system performance [14]. On one hand, as analyzed in Sec. II-A, the rank of the channel matrix is directly related to $K$, thus affecting the SDoF and the sum SE. On the other hand, the values of $K$ and $d_{s}$ affect the position of each antenna, and the channel matrix will change accordingly, resulting in better or worse sub-channels. In practice, due to the space constraint, the size of the BS array aperture must be smaller than a certain threshold, e.g., between $0.2$m and $1.2$[14, 33]. Therefore, the optimization problem aims to jointly design the $(K,d_{s})$ pair along with the hybrid precoders and combiners to maximize the sum SE. Based on (4) and the constraints, the overall optimization problem for FC can be formulated as
@@
$S_{t}^{\text{min}}\leq S_{t}\leq S_{t}^{\text{max}},$ (7e)
where constraint (7e) denotes the array aperture constraint where $S_{t}^{\text{min}}$ refers to the case of a compact array. While for SC, the constraint (7c) is replaced by $|\mathbf{F}_{j}(m,n)|^{2}=\frac{1}{N_{t}/K}$. This is a non-convex problem due to the constant-modulus constraints of the analog beamformers. Moreover, the effect of $(K,d_{s})$ is implicit within the expression of the channel matrix, which further complicates the design problem, making it intractable to solve directly. Therefore, we separate this problem into two sub-problems, for which we first design the MU-WSA architecture by analyzing the impact of the number of subarrays and the subarray spacing. Then, we design the hybrid beamforming algorithm to maximize the sum SE.
@@ III Design of MU-WSA Architecture
In this section, we focus on the design of the architecture and neglect the influence of the beamforming algorithm, in order to analyze the upper bound of the sum SE. That is to say, we assume each user transmits its own data streams without interference, such that the performance of each user can be analyzed independently. Considering equal power allocation, the optimal SE, i.e., capacity, for each user is written as
@@
where $\rho=P_{t}/\operatorname{rank(\mathbf{H}_{u})}$ denotes the allocated power for each data stream. Since the THz channel has the line-of-sight (LoS)-dominant feature [34], we further simplify the optimization problem as to maximize the capacity of the LoS channel. Moreover, since the correlation of the array response vectors with different elevation angles and different azimuth angles are approximately zero when the number of antennas approaches infinity [21, 35], the rank of the LoS channel approximately equals $K$. To this end, the design problem for the MU-WSA architecture is formed as
@@
$\bar{\mathbf{A}}_{r}(i,j)$ $=\sum_{k=1}^{K}e^{j\frac{2\pi}{\lambda}d_{a}[(j_{1}-i_{1})\sin(\theta_{t}^{uk})\cos(\phi_{t}^{uk})+(j_{2}-i_{2})\sin(\phi_{t}^{uk})]},$ (14a)
@@
Since f increases monotonically with $|\bar{\mathbf{A}}_{r}(i,j)|^{2}$, we can further transform the minimization of f as the minimization of function $|\bar{\mathbf{A}}_{r}(i,j)|^{2},\forall i,j$. Define $\bar{\mathbf{A}}_{r}(i,j)=\sum_{k=1}^{K}e^{j\beta_{k,i,j}}$. Since the azimuth angle and the elevation angle can be calculated based on the geometry shown in Fig. 1, i.e.,
$\cos(\phi_{t}^{uk})$ $=\frac{z_{u}-z_{k}}{D^{uk}},$ (16a)
$\sin(\theta_{t}^{uk})$ $=\frac{x_{u}-x_{k}}{D^{uk}\cos{\phi_{t}^{uk}}},$ (16b)
then $\beta_{k,i,j}$ can be transferred into a function of $d_{s}$. Therefore, we have
@@
Remark 1: In the proof, we impose a stricter condition than necessary by assuming that each summand $|\mathbf{A}_{r}(i,j)|$ in the function $f$ is monotonically increasing. As a result, the threshold of the transmission distance, i.e., $\tau=2\sqrt{2}(\sqrt{N_{r}}-1)S_{t}^{\text{max}}<D^{uk}$, is in fact a upper bound of the real threshold. That is to say, in practice, the distance can be smaller than $\tau$, while the monotonicity still holds.
@@
To directly illustrate the relationship between the channel capacity and the subarray spacing, we present the numerical analysis in Fig. 2. The number of antennas at the BS and each user is set as $N_{t}=1024$ and $N_{r}=64$. Since $K$ denotes the number of subarrays, it must be a factor of the number of antennas. What’s more, for simplicity, we assume $K$ is a square. Therefore, we exhaustively sweep the choices of $K$ in $\mathcal{K}=[1,4,16,64,256]$. For each possible value of $K$, we vary the subarray spacing varies from $1-1000\lambda$. Since the array aperture cannot be infinitely large, we omit the case for larger subarray spacing. In the figure, different colored lines represent different values of $K$, and each line shows the relationship between the capacity and the subarray spacing. At $K=1$, the horizontal line is the baseline representing the compact array case, where the capacity is invariable with the change of the subarray spacing. While for other values of $K$, the channel capacity increases monotonically with the subarray spacing. This is because when $d_{s}$ increases, the difference of the angles $\phi_{p}^{(uk)}$ and $\theta_{p}^{(uk)}$ in (1) becomes larger, so that the difference between the sub-matrices $\mathbf{H}^{(uk)}$ is also larger, leading to a better spatial multiplexing gain. Therefore, based on Theorem 1, and the fact that the choices of $K$ are limited, we propose to first calculate the maximum subarray spacing $d_{s}^{K,\text{max}}$ for each possible value of $K$ under the array aperture threshold, which can be calculated as
$d_{s}^{K,\text{max}}=\frac{\sqrt{2}S_{t}^{\text{max}}+\lambda(\sqrt{N_{t}}-\sqrt{K})}{2(\sqrt{K}-1)}.$ (19)
Then we exhaustively search from all the possible $(K,d_{s})$ pairs to find the solution with the highest channel capacity. This procedure is summarized in Algorithm 1.
Algorithm 1 Search-based design algorithm for subarray sapcing and subarray number in MU-WSA architecture.
0:  Channel matrix $\mathbf{H}$, array aperture threshold $S_{t}^{\text{max}}$.
1:  for $k\in\mathcal{K}$ do
2:     Calculate the maximum available $d_{s}$ through (19) such that the array aperture is no larger than $S_{t}^{\text{max}}$.
3:  end for
4:  Select $(K,d_{s})$ with the highest sum SE.
@@ IV Design of Hybrid Beamforming Algorithm
In this section, we aim to design the hybrid beamformers to maximize the sum SE for all users. To simplify the overall beamforming design problem, we separately design the digital and the analog stages. For the digital beamformers, the BD algorithm is applied to eliminate user interference. For the design of the analog beamformer, two different connection types are considered, i.e., the sub-connected and fully-connected WSA. For the sub-connected case, we propose an alternating-optimization-based algorithm, while for the FC architecture, we propose two low-complexity subarray-based algorithms.
@@ IV-A Digital Precoder and Combiner
where $\hat{\mathbf{V}}_{u}^{(1)}$ is the first $(U-1)L_{r}$ right singular vectors of $\hat{\mathbf{H}}_{u}$, and $\hat{\mathbf{V}}_{u}^{(0)}$ denotes the orthogonal bases of the null space of $\hat{\mathbf{H}}_{u}$, i.e., $\bar{\mathbf{H}}_{u}\hat{\mathbf{V}}_{i}^{(0)}=0,i\neq u$. As a result, we can eliminate the inter-user interference so that each user can transmit data stream via their own sub-channel $\bar{\mathbf{H}}_{u}\hat{\mathbf{V}}_{u}^{(0)}$. Without the inter-user interference, the sum SE (7a) can then be maximized by achieving the maximum transmission SE of each user. Therefore, we further optimize the precoder and combiner on each sub-channel via SVD, i.e., $\bar{\mathbf{H}}_{u}\hat{\mathbf{V}}_{u}^{(0)}=\mathbf{U}_{u}\mathbf{\Sigma}_{u}\mathbf{V}_{u}^{H}$. We denote $\mathbf{F}_{\mathrm{BB}}=\left[\mathbf{F}_{\mathrm{BB},1},\cdots,\mathbf{F}_{\mathrm{BB},U}\right]$, where $\mathbf{F}_{\mathrm{BB},u}$ is the baseband precoder for the $u^{\textrm{th}}$ data stream. The expressions of the baseband precoder and combiner are given by
$\mathbf{F}_{\mathrm{BB}}$ $=\left[\mathbf{F}_{\mathrm{BB},1},\cdots,\mathbf{F}_{\mathrm{BB},U}\right]$
$=\left[\hat{\mathbf{V}}_{1}^{(0)}\mathbf{V}_{1}^{\left(N_{s}\right)},\cdots,\hat{\mathbf{V}}_{U}^{(0)}\mathbf{V}_{U}^{\left(N_{s}\right)}\right],$ (21a)
@@ IV-B Analog Precoder and Combiner
As in UM-MIMO systems, the optimal analog combiners typically satisfy $\mathbf{W}_{\mathrm{RF}}^{H}\mathbf{W}_{\mathrm{RF}}\approx\mathbf{I}$ [36], the objective function can be simplified as
$SE\approx\log_{2}\operatorname{det}\left(\mathbf{I}+\sigma_{\mathrm{n}}^{-2}\mathbf{W}_{\mathrm{RF}}^{H}\mathbf{H}\mathbf{F}_{\mathrm{RF}}\mathbf{F}_{\mathrm{RF}}^{H}\mathbf{H}^{H}\mathbf{W}_{\mathrm{RF}}\right).$ (23)
IV-B1 Alternating-Optimization-based Analog Beamforing in Sub-connected MU-WSA
In the sub-connected MU-WSA system where the special block-diagonal structure of the analog precoder should be considered as in (6). To solve this problem, we propose to isolate the contribution of each sub-matrix and optimize one sub-matrix per iteration with other sub-matrices fixed.
We first consider the design of the analog precoder assuming fixed $\mathbf{W}_{\mathrm{RF}}$. Then, the matrix in (23) can be decomposed into four parts, i.e.,

### v2

@@ II-B2 Sub-connected MU-WSA
In the sub-connected WSA, we consider a more general scenario where each subarray is connected to multiple RF chains, in contrast to traditional SC architectures that assign one RF chain per subarray as stated in Sec. I. In this context, the $L_{t}$ RF chains are divided uniformly into $K$ groups, with each group connected to one subarray. This results in a reduction of phase shifters by a factor of $K^{2}$ compared to FC systems. As a result, the analog precoder adopts a block-diagonal structure as
$\mathbf{F}_{\mathrm{RF}}=\begin{bmatrix}\mathbf{F}_{1}&0&\ldots&0\\ 0&\mathbf{F}_{2}&\ldots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&\ldots&0&\mathbf{F}_{K}\end{bmatrix},$ (7)
where $\mathbf{F}_{j}\in\mathbb{C}^{\frac{N_{t}}{k}\times\frac{L_{t}}{k}}$ represents the matrix of the analog precoder for each subarray, and should satisfy $\left|\mathbf{F}_{j}(m,n)\right|=\frac{1}{\sqrt{N_{t}/K}},\forall j$.
@@ II-C Problem Formulation
In MU-WSA architecture, the subarray spacing $d_{s}$ and the number of subarrays $K$ need to be investigated since they have a significant impact on the system performance [16]. On one hand, as analyzed in Sec. II-A, the rank of the channel matrix is directly related to $K$, thus affecting the SDoF and the sum SE. On the other hand, the values of $K$ and $d_{s}$ affect the positions of each antenna and each subarray, which further influence the transmission distance $D_{p}^{uk}$ and the array response vectors $\mathbf{a}_{tp}^{uk}$ and $\mathbf{a}_{rp}^{uk}$ of the channel matrix in eq. (3).
Specifically, by denoting the bottom right antenna as the reference antenna of each subarray, the position of the first subarray can be written as $p_{1}=(0,0,h_{t})$, where $h_{t}$ denotes the height of the BS. For simplicity, we assume the number of antennas in each subarray on the x- and z-axis are the same, which equals $n=\sqrt{\frac{N_{t}}{K}}$. Therefore, the distance between different subarrays on the x- and z-axis are also the same, which can be calculated as $d=\frac{(n-1)}{2}\lambda+d_{s}$. We denote the position of the $k^{\mathrm{th}}$ subarray as $p_{k}=(x_{k},0,z_{k})$. Then, based on the geometry shown in Fig. 1, we can derive
$x_{k}$ $=x_{1}+(k_{x}-1)d=(k_{x}-1)(\frac{(n-1)}{2}\lambda+d_{s}),$ (8)
$z_{k}$ $=z_{1}+(k_{z}-1)d=h_{t}+(k_{z}-1)(\frac{(n-1)}{2}\lambda+d_{s}),$
where $k_{x}=k-(k_{z}-1)\sqrt{K}$ and $k_{z}=\lceil\frac{k}{\sqrt{K}}\rceil$ are the index positions of the $k^{\mathrm{th}}$ subarray on the x- and z-axis, respectively. Then, the transmission distance of the LoS path can be calculated as
$D_{1}^{uk}=\sqrt{(x_{k}-x_{u})^{2}+(y_{k}-y_{u})^{2}+(z_{k}-z_{u})^{2}}$ (9)
$=\sqrt{((k_{x}-1)d-x_{u})^{2}+y_{u}^{2}+(h_{t}+(k_{z}-1)d-z_{u})^{2}},$
where $p_{u}=(x_{u},y_{u},z_{u})$ is the position of user $u$. Moreover, the azimuth and elevation angles for the LoS path are also influenced by $(K,d_{s})$ as
$\cos(\phi_{t}^{uk})$ $=\frac{z_{u}-z_{k}}{D_{1}^{uk}}=\frac{z_{u}-h_{t}-(k_{z}-1)d}{D_{1}^{uk}},$ (10a)
$\sin(\theta_{t}^{uk})$ $=\frac{x_{u}-x_{k}}{D_{1}^{uk}\cos{\phi_{t}^{uk}}}=\frac{x_{u}-(k_{x}-1)d}{D_{1}^{uk}\cos{\phi_{t}^{uk}}}.$ (10b)
For NLoS paths, the angles and the transmission distance can be calculated similarly.
As the $(K,d_{s})$ pair influences the channel matrix and affects the system performance, it requires careful optimization. However, due to the limited space in practice, the size of the array aperture at the BS is constrained to be smaller than a certain threshold, e.g., between $0.2$ m and $1.2$ m [16, 40]. Considering a square UPA with array aperture $S_{t}$ representing the length of its diagonal, we have
$S_{t}=(\sqrt{\frac{N_{t}}{2}}-\sqrt{\frac{K}{2}})\lambda+(\sqrt{2K}-\sqrt{2})d_{s}.$ (11)
Therefore, the optimization of $(K,d_{s})$ is constrained by the maximum array aperture such that they can not be arbitrarily large or small.
Apart from the design of $(K,d_{s})$ pair, the hybrid beamformers also need to be optimized to maximize the sum SE for both FC and SC systems. In FC systems, the overall optimization problem can be formulated as
@@
$S_{t}^{\text{min}}\leq S_{t}\leq S_{t}^{\text{max}}.$ (12e)
Constraint (12e) denotes the array aperture constraint where $S_{t}^{\text{min}}$ refers to the case of a compact array. For SC, constraint (12c) is replaced by $|\mathbf{F}_{j}(m,n)|^{2}=\frac{1}{N_{t}/K}$. This is a non-convex problem due to the constant-modulus constraints of the analog beamformers. Moreover, the effect of $(K,d_{s})$ is implicit within the expression of the channel matrix, which further complicates the design problem and makes it intractable to solve directly. Therefore, we separate this problem into two sub-problems, for which we first design the MU-WSA architecture by analyzing the impact of the number of subarrays and the subarray spacing. After that, the channel matrix can be determined, based on which we design the hybrid beamformers to maximize the sum SE as the second sub-problem.
@@ III Design of MU-WSA Architecture
In this section, we focus on the optimization of the MU-WSA architecture, i.e., design of the $(K,d_{s})$ pair, to maximize the upper bound of the sum SE, which can be achieved by assuming optimal precoding and combining [10, 41]. That is to say, we assume each user transmits its own data streams without interference, such that the performance of each user can be analyzed independently. Considering equal power allocation, the optimal SE, i.e., capacity, for each user is written as
@@
where $\rho=P_{t}/\operatorname{rank(\mathbf{H}_{u})}$ denotes the allocated power for each data stream, and $r_{i}(\mathbf{H}_{u})$ represents the $i^{\mathrm{th}}$ singular value of $\mathbf{H}_{u}$. Since the THz channel has the LoS-dominant feature [16], we further simplify the optimization problem as to maximize the capacity of the LoS channel, i.e., $N_{p}=1$. Based on (3), the LoS channel can be derived as
$\mathbf{H}_{u}^{LoS}=\begin{bmatrix}\mathbf{H}^{u1}&\cdots&\mathbf{H}^{uk}\end{bmatrix}$ (14)
[…]
Moreover, since the correlation of the array response vectors with different elevation angles and different azimuth angles are approximately zero when the number of antennas approaches infinity [23, 42], the rank of the LoS channel approximately equals $K$. To this end, the design problem for the MU-WSA architecture is formed as
@@
$\bar{\mathbf{A}}_{r}(i,j)=\sum_{k=1}^{K}e^{j\frac{2\pi}{\lambda}d_{a}[(j_{1}-i_{1})\sin(\theta_{t}^{uk})\cos(\phi_{t}^{uk})+(j_{2}-i_{2})\sin(\phi_{t}^{uk})]},$ (20)
@@
Since f increases monotonically with $|\bar{\mathbf{A}}_{r}(i,j)|^{2}$, we can further transform the minimization of f as the minimization of function $|\bar{\mathbf{A}}_{r}(i,j)|^{2},\forall i,j$. Define $\bar{\mathbf{A}}_{r}(i,j)=\sum_{k=1}^{K}e^{j\beta_{k,i,j}}$. Then, based on eq. (10), $\beta_{k,i,j}$ can be transferred into a function of $d_{s}$. Therefore, we have
@@
Remark 1: In the proof, we impose a stricter condition than necessary by assuming that each summand $|\mathbf{A}_{r}(i,j)|$ in the function $f$ is monotonically increasing. As a result, the threshold of the transmission distance, i.e., $\tau=2\sqrt{2}(\sqrt{N_{r}}-1)S_{t}^{\text{max}}$, is in fact a upper bound of the real threshold. That is to say, in practice, the distance can be smaller than $\tau$, while the monotonicity still holds.
@@
Remark 3: The prerequisite of the Theorem, i.e., $\tau=2\sqrt{2}(\sqrt{N_{r}}-1)S_{t}^{\text{max}}<D^{uk}$, imposes a constraint that the array aperture must not be too large compared to the transmission distance. Therefore, in DAS where the distance between different subarrays can achieve tens of meters such that the array aperture and the transmission distance are often comparable, this Theorem is not applicable.
@@
To directly illustrate the relationship between the channel capacity and the subarray spacing, we present the numerical analysis in Fig. 2. The numbers of antennas at the BS and each user are set as $N_{t}=1024$ and $N_{r}=16$. We exhaustively sweep the choices of $K$ in $\mathcal{K}=[1,4,16,64,256]$. For each possible value of $K$, we vary the subarray spacing varies from $1-8000\lambda$. Since the array aperture cannot be infinitely large, we omit the case for larger subarray spacing. In the figure, different colored lines represent different values of $K$, and each line shows the relationship between the capacity and the subarray spacing. At $K=1$, the horizontal line is the baseline representing the compact array case, where the capacity is invariable with the change of the subarray spacing. For other values of $K$, the channel capacity first increases monotonically with the subarray spacing. This is because when $d_{s}$ increases, the difference of the angles $\phi_{p}^{(uk)}$ and $\theta_{p}^{(uk)}$ in (1) becomes larger, so that the difference between the sub-matrices $\mathbf{H}^{(uk)}$ is also larger, leading to a better spatial multiplexing gain. While as the subarray spacing continues to increase, eventually the prerequisite in Theorem 1 no longer holds, and the capacity begins to decrease. Note that when the number of subarrays is relatively small, i.e., $K=4$, the array aperture changes slowly with respect to the subarray spacing. As a result, the decrease in capacity is not immediately obvious. However, upon closer inspection of the figure, it is clear that the capacity still shows a declining trend from $40.6$ bps/Hz at $d_{s}=6400\lambda$ to $40.3$ bps/Hz at $d_{s}=8000\lambda$. This occurs because, when the array aperture becomes comparable to the transmission distance, the subarray spacing starts to have a more significant impact on the transmission distance. In other words, a larger subarray spacing leads to a larger transmission distance, which results in higher path loss and reduced channel capacity. However, in our MU-WSA systems, the subarray spacing is typically much smaller than the point at which the capacity starts to decrease. As a result, we can safely conclude that, in MU-WSA systems, the channel capacity increases with larger subarray spacing.
This monotonicity is particularly suitable to a multi-user scenario. As the channel environment is typically dynamic, it is impractical to constantly adjust the positions of subarrays. Fortunately, based on the previous analysis, a positive conclusion can be drawn that as long as the number of subarrays and the array aperture are determined, the subarray spacing only needs to be set to the maximum, regardless of how the users change their positions. In practice, the number of subarrays and the array aperture are affected and determined by hardware constraint. For example, the number of subarrays should be smaller than the number of RF chains, and the available size for the antenna array is usually limited. Therefore, with fixed number of subarrays and maximum array aperture, the optimal subarray spacing can be easily obtained through
$d_{s}^{K,\text{max}}=\frac{\sqrt{2}S_{t}^{\text{max}}-\lambda(\sqrt{N_{t}}-\sqrt{K})}{2(\sqrt{K}-1)}.$ (24)
In this way, real-time adjustments of the architecture is unnecessary.
@@ IV Design of Hybrid Beamforming Algorithm
After optimizing the array structure, the channel matrix can be determined for the design of the beamformers. To further simplify the beamforming sub-problem, we separately consider the digital and the analog stages. For the digital beamformers, the BD algorithm is applied to eliminate user interference. For the design of the analog beamformer, two different connection types are considered, i.e., the sub-connected and fully-connected WSA. For the sub-connected case, we propose an alternating-optimization-based algorithm, while for the FC architecture, we propose two low-complexity subarray-based algorithms.
@@ IV-A Digital Precoder and Combiner
where $\hat{\mathbf{V}}_{u}^{(1)}$ is the first $(U-1)L_{r}$ right singular vectors of $\hat{\mathbf{H}}_{u}$, and $\hat{\mathbf{V}}_{u}^{(0)}$ denotes the orthogonal bases of the null space of $\hat{\mathbf{H}}_{u}$, i.e., $\bar{\mathbf{H}}_{u}\hat{\mathbf{V}}_{i}^{(0)}=0,i\neq u$. As a result, we can eliminate the inter-user interference so that each user can transmit data stream via their own sub-channel $\bar{\mathbf{H}}_{u}\hat{\mathbf{V}}_{u}^{(0)}$. Without the inter-user interference, the sum SE (12a) can then be maximized by achieving the maximum transmission SE of each user. Therefore, we further optimize the precoder and combiner on each sub-channel via SVD, i.e., $\bar{\mathbf{H}}_{u}\hat{\mathbf{V}}_{u}^{(0)}=\mathbf{U}_{u}\mathbf{\Sigma}_{u}\mathbf{V}_{u}^{H}$. Then, the expressions of the baseband precoder and combiner are given by
$\mathbf{F}_{\mathrm{BB}}=\left[\hat{\mathbf{V}}_{1}^{(0)}\mathbf{V}_{1}^{\left(N_{s}\right)},\cdots,\hat{\mathbf{V}}_{U}^{(0)}\mathbf{V}_{U}^{\left(N_{s}\right)}\right],$ (26a)
@@ IV-B Analog Precoder and Combiner
As in UM-MIMO systems, the optimal analog combiners typically satisfy $\mathbf{W}_{\mathrm{RF}}^{H}\mathbf{W}_{\mathrm{RF}}\approx\mathbf{I}$ [43, 7], the objective function can be simplified as
$SE$ $\approx\log_{2}\operatorname{det}\left(\mathbf{I}+\sigma_{\mathrm{n}}^{-2}\mathbf{W}_{\mathrm{RF}}^{H}\mathbf{H}\mathbf{F}_{\mathrm{RF}}\mathbf{F}_{\mathrm{RF}}^{H}\mathbf{H}^{H}\mathbf{W}_{\mathrm{RF}}\right)$ (28a)
$=\log_{2}\operatorname{det}\left(\mathbf{I}+\sigma_{\mathrm{n}}^{-2}\mathbf{F}_{\mathrm{RF}}^{H}\mathbf{H}^{H}\mathbf{W}_{\mathrm{RF}}\mathbf{W}_{\mathrm{RF}}^{H}\mathbf{H}\mathbf{F}_{\mathrm{RF}}\right).$ (28b)
IV-B1 Alternating-Optimization-based Analog Beamforming in Sub-connected MU-WSA
In the sub-connected MU-WSA system where the special block-diagonal structure of the analog precoder should be considered as in (7). Since we consider a general case where each subarray can be connected to more than one RF chain, each block in $\mathbf{F}_{\mathrm{RF}}$ can be a matrix. To solve this problem, we propose to isolate the contribution of each sub-matrix and optimize one sub-matrix per iteration with other sub-matrices fixed.
We first consider the design of the analog precoder assuming fixed $\mathbf{W}_{\mathrm{RF}}$. Then, the matrix in (28b) can be decomposed into four parts, i.e.,